A rocket rising from the ground has a velocity of
v(t) = 2,000 te−t/180 ft/s,
after t seconds. How far does it rise in the first 3 minutes? (Round your answer to the nearest foot.)
1
Expert's answer
2019-11-18T13:39:34-0500

v(t)=2000te-t/180; "\," v(180)=2000*180e-180/180=132437 ft